OpenStudy (anonymous):

A bullet of mass 20g moving with a speed of 150m/s atrikes a block and comes to rest after piercing 10cm into it . Find the avg. force of resistance offered by the block.

OpenStudy (anonymous):

@akash123

OpenStudy (anonymous):

find out the change in momnetum

OpenStudy (anonymous):

then average force of resistance= rate of change of momentum= (change in momentum)/ (time interval)

OpenStudy (anonymous):

so 1st u have to calculate..how much time bullet will take in coming to rest?

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

i guess i dont understand , sorry akash am slow in this , and i want you to help me , how to calculate change in momentum ?

OpenStudy (anonymous):

change in momentum= (final momentum- initial momentum)

OpenStudy (anonymous):

1st find accl

OpenStudy (anonymous):

i mean... 1st the bullet was moving with a speed of 150m/s then its momentum = m(150), m=20g

OpenStudy (anonymous):

yes use the formula v^2- u^2= 2as

OpenStudy (anonymous):

\[V^2 - u^2 = 2as\] \[\frac{ 150*150*100 }{ 2*10 } = a\]

OpenStudy (anonymous):

\[F= Ma\]

OpenStudy (anonymous):

if u find the acceleration, then u can calculate the average resistance force by multiplying acceleration with mass

OpenStudy (anonymous):

Note mass shuld be in kg 1 gm = 10^-3 kg

OpenStudy (anonymous):

no need to calculate change in the momentum

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

According to COD ...this is maths page...and u r asking physics...Question....Plzzz...Go..to physics page

OpenStudy (anonymous):

okay let me try , @akash123 and @Yahoo! thank yu !

OpenStudy (anonymous):

am sry i din see that .. am sry !

OpenStudy (anonymous):

No...Problem.....and welcomee...

OpenStudy (anonymous):

:)